# Video LLMs for Temporal Reasoning in Long Videos

### v1

@@ Abstract
This paper introduces TemporalVLM, a video large language model capable of effective temporal reasoning and fine-grained understanding in long videos. At the core, our approach includes a visual encoder for mapping a long-term input video into features which are time-aware and contain both local and global cues. In particular, it first divides the input video into short-term clips, which are jointly encoded with their timestamps into time-sensitive local features. Next, the local features are passed through a bidirectional long short-term memory module for global feature aggregation. The extracted time-aware and multi-level features are important for accurate temporal reasoning and fine-grained understanding in long videos. Moreover, to facilitate the evaluation of TemporalVLM, we present a large-scale long video dataset of industry assembly processes, namely IndustryASM, which consists of videos recorded on factory floors with actions and timestamps annotated by industrial engineers for time and motion studies and temporal action segmentation evaluation. Finally, extensive experiments on datasets of long videos, including TimeIT and IndustryASM, show that TemporalVLM achieves superior performance than previous methods across temporal reasoning and fine-grained understanding tasks, namely dense video captioning, temporal video grounding, video highlight detection, and temporal action segmentation.
@@ 1 Introduction
Video temporal reasoning represents the process of reasoning about time and its passage in videos, with a focus on how events or actions happen and relate to each other in terms of time, e.g., when the actions or events happen, their durations, their orders, and their causal relationships in the videos. Several video understanding applications require the ability of temporal reasoning. Examples include dense video captioning [51, 64, 58], which attempts to detect the timestamps and generate detailed descriptions of all the events or actions in an input video, and temporal action segmentation [28, 59, 2, 30], which aims to associate each frame of a long video capturing a complex activity with one of the action/sub-activity classes. Considerable efforts [58, 31, 34, 30] have been invested in developing separate models for solving individual tasks. These models usually have different architectures. Therefore, it is favorable to design a unified model for handling various tasks.
@@
In this paper, we propose TemporalVLM, a video LLM capable of effective temporal reasoning and fine-grained understanding in long videos. Fig. 1 shows the differences in model architectures of TemporalVLM and previous works. Our model includes two key architectural contributions for temporal reasoning in long videos. Firstly, we divide a long-term input video into multiple short-term clips and introduce a time-sensitive clip encoder for extracting time-aware local features from each clip. Secondly, we adopt a bidirectional long short-term memory (BiLSTM) module [12] which takes all the local features as inputs and computes global features from multiple clips. As a consequence, our features not only are time-sensitive but also contain both local fine-grained and global semantic information, which are crucial for temporal reasoning in long videos. Moreover, to further evaluate the performance of our model, we present IndustryASM, a large-scale long video dataset of industry assembly processes for temporal action segmentation benchmarking [28, 59, 24, 22, 2, 49, 19, 30] and time and motion studies. Our IndustryASM dataset comprises of $4803$ videos with an average video duration of $119$ seconds. It covers in total $52$ diverse industry assembly tasks and includes timestamp and action labels. We convert the labels into chat samples with manually written instructions. To the best of our knowledge, it is the first large-scale video dataset recorded on factory floors and labeled by industrial engineers for temporal action segmentation evaluation and time and motion analyses. Lastly, extensive experiments on datasets of long videos, i.e., TimeIT [40] and IndustryASM, demonstrate that our TemporalVLM approach achieves superior results over previous methods across temporal reasoning and fine-grained understanding tasks, namely dense video captioning, temporal video grounding, video highlight detection, and temporal action segmentation.
@@
  • We develop TemporalVLM, a video LLM for temporal reasoning and fine-grained understanding in long videos.
  • Given a long-term input video, we first sample short-term clips and encode each clip along with its timestamps into time-aware local features. The local features are then fed to a BiLSTM module to extract global features. Therefore, our features not only are time-sensitive but also capture effective local and global information.
@@
  • Our evaluations show that TemporalVLM outperforms prior works across temporal reasoning and fine-grained understanding tasks on datasets of long videos.
@@ 2 Related Work
Video Large Language Models. Motivated by the impressive reasoning and generalization abilities of LLMs, video LLMs [27, 29, 32, 33, 61, 20, 43] have been developed for video understanding tasks. They typically include three key components: i) a pre-trained visual encoder (e.g., ViT [42]) to extract visual features, ii) a projection layer to map visual features into the text latent space of LLMs, and iii) a pre-trained LLM (e.g., LLaMA [46]) for generating responses. They mostly differ in their visual encoder, i.e., how visual features are extracted and aggregated from an input video. For example, VideoChat [27] extracts frame features via a visual transformer [42] and employs a query transformer (Q-Former) [26] to aggregate frame features into video features, while a video Q-Former is further included for temporal modeling in Video-LLaMA [61]. The above methods usually map the video into a fixed number of visual tokens, yielding degrading performance with long videos, while encoding video frames and their timestamps separately and hence struggling with temporal reasoning tasks. To overcome the above limitations, a few methods with temporal reasoning capabilities have been introduced, e.g., TimeChat [40] presents a sliding video Q-Former to handle various video lengths and a time-aware frame encoder to jointly encode video frames and their timestamps. These methods often consider the video as a single clip [32, 33, 27, 61] and aggregate video tokens via pooling operation [32, 33] and query aggregation [27, 61], overlooking fine-grained details. In this work, we divide the video into multiple clips and propose a time-aware clip encoder for capturing local features from each clip. In addition, we introduce a BiLSTM module for aggregating global features over multiple clips.
@@
Long Video Understanding. There exist a few challenges in long video understanding such as complex spatial-temporal relationships and redundant information. Several methods have been developed for long video understanding, including efficient architectures [8, 23], temporal pooling/aggregation [41, 56], and dynamic clip selection [11, 13]. For vision-language understanding tasks, methods based on temporal alignment [15, 14, 3, 48] have been introduced, while memory techniques [57, 62] are often used in dense video prediction tasks. Recently, Weng et al. [55] propose LongVLM for long video understanding. In particular, it divides the video into clips and employs a merging module for extracting local features and pooling operation for computing global features from multiple clips. It does not utilize timestamp information, which is crucial for temporal reasoning. In contrast, our TemporalVLM model explicitly utilizes timestamps via a time-aware clip encoder, yielding time-aware features for temporal reasoning. Moreover, we employ a learnable BiLSTM module for aggregating global features from multiple clips.
@@ 3 TemporalVLM: A Video LLM for Temporal Reasoning in Long Videos
Figure 2: Given a long-term input video with timestamps, we first divide it into short-term clips with timestamps. The time-aware clip encoder is then employed on each clip along with its timestamps for extracting time-aware local features, which capture local fine-grained cues within each clip. Next, the bidirectional long short-term memory module is adopted for computing global features, which capture global temporal relationships across multiple clips. Lastly, the large language model takes both video and text tokens as inputs and generates responses to users.
In the following, we describe our main contribution, namely TemporalVLM — a video LLM for effective temporal reasoning in long videos. Two novel components are proposed in TemporalVLM: i) a time-aware clip encoder for extracting time-aware local features from a short-term clip, and ii) a BiLSTM module for computing global features from the local features.
3.1 Overall Architecture
We provide an overview of our TemporalVLM’s architecture in Fig. 2, including three key components: a time-aware clip encoder, a BiLSTM module, and a LLM. The time-aware clip encoder maps a short-term clip sampled from the long-term input video into time-aware local features. Next, the BiLSTM module converts local features into global features. Finally, the LLM takes the extracted features along with instructions as inputs and generates responses.
@@ 3.2 Time-Aware Clip Encoder
Existing video LLMs usually consider a long-term input video as a single clip and uniformly sample a fixed number of frames $N^{v}_{f}$ over the entire video duration. For instance, TimeChat [40] samples $N^{v}_{f}=96$ frames from a given input video, while VTimeLLM [17] samples $N^{v}_{f}=100$ frames. In addition, previous works usually employ query aggregation, e.g., TimeChat [40], and pooling operation, e.g., Video-ChatGPT [33], to aggregate video tokens. As a consequence, they often perform well with short-term input videos but miss out important fine-grained information for long-term input videos, leading to degrading performance.
To alleviate the above issue, we propose a time-aware clip encoder (i.e., blue block in Fig. 2) for extracting local fine-grained cues within a clip. Given a long-term input video, we first decompose it into $C$ short-term clips and uniformly sample a fixed number of frames $N^{c}_{f}$ from each clip. Here, we set $C=6$ clips and $N^{c}_{f}=96$ frames per clip. The sampled frames for each clip along with their corresponding timestamps are then passed to our time-aware clip encoder for simultaneously encoding the frame contents and their timestamps, yielding time-aware local features. Specifically, we adopt the time-aware frame encoder and sliding video Q-Former of TimeChat [40] as our time-aware clip encoder. The time-aware frame encoder uses a pre-trained image encoder [44] to obtain frame features, which are jointly encoded with timestamps via an image Q-Former [7], yielding time-aware frame features. These features are then fed to a sliding video Q-Former [40] for temporal modeling, yielding time-aware clip features for each clip. A few recent works, e.g., LongVLM [55], leverage clip sampler and clip encoder for long video understanding. However, their encoders are not time-aware and hence are not effective for temporal reasoning in long videos.
@@ 3.3 Bidirectional Long Short-Term Memory
The time-sensitive clip encoder is applied on each short-term clip to obtain time-aware local features, which capture useful fine-grained cues for temporal reasoning within the short-term clip. However, they are not effective for temporal reasoning over the long-term video, which requires the ability to capture long-range temporal dependencies.
In this section, we propose a BiLSTM module [12] (i.e., green block in Fig. 2) for computing global features across multiple clips. Particularly, we first concatenate time-aware local features extracted from clips in the temporal order which the clips appear in the video. We then pass the sequence of local features to our BiLSTM module for aggregating global features. As we will show later in Sec. 5.3, a BiLSTM module, which is capable of capturing long-range temporal relationships in both directions, leads to better performance than a traditional LSTM module, which conducts the forward pass only. As a result, our features not only are time-sensitive but also contain both local fine-grained and global semantic information, which are important for temporal reasoning in long videos. Unlike the recent LongVLM model [55], which applies pooling operation to obtain global features, our BiLSTM module has learnable parameters and yields better results as demonstrated in Sec. 5.
The video tokens output by our BiLSTM module have dimensions of $(C*N^{c}_{f},~{}2*N_{V})$, while the LLM requires an input of dimensions $(N^{c}_{f},~{}N_{LLM})$. Here, $N_{V}$ and $N_{LLM}$ are the dimensions of the video tokens and the LLM latent space respectively, and $2$ is to account for the outputs of both forward and backward passes of BiLSTM. Therefore, we pass the video tokens output by BiLSTM through projection layers to match the input dimensions required by the LLM.
@@ 3.4 Large Language Model
The LLM (i.e., yellow block in Fig. 2) takes as input the video tokens $\bm{X}^{v}$, query tokens $\bm{X}^{q}$, and generates responses $\bm{X}^{r}$ to users. Video LLMs typically employ a two-stage training scheme: i) leveraging large-scale image/video-text pairs for vision-language alignment to pre-train the model, and ii) utilizing instruction data for instruction following to fine-tune the pre-trained model. Similar to TimeChat [40], we use the checkpoint of the open-source LLaMA-2 7B model [47] and perform instruction tuning only. We optimize the below loss function for language modeling:
$\mathcal{L}=-\log P_{\bm{\psi}}(\bm{X}^{r}|\bm{X}^{v},\bm{X}^{q})$ (1)
$=-\sum^{L}_{i=1}\log P_{\bm{\psi}}(x_{i}|\bm{X}^{v},\bm{X}^{q},\bm{X}^{r}_{<i}).$ (2)
In the above, $\bm{\psi}$ is the learnable parameters of our TemporalVLM model, $L$ denotes the response length, $x_{i}$ is the current predicted token, and $\bm{X}^{r}_{<i}$ represents the previous tokens appearing before $x_{i}$ in the response. In addition, we employ LoRA [16] for efficient fine-tuning.
@@ 4 IndustryASM: A Large-Scale Long Video Dataset of Industry Assembly Processes
We follow prior works [37, 17, 18, 40] to test our TemporalVLM model on three temporal reasoning tasks, namely dense video captioning, temporal video grounding, and video highlight detection. To further evaluate the temporal reasoning performance of our method, we introduce IndustryASM, a large-scale long video dataset of manufacturing assembly procedures for temporal action segmentation benchmarking and time and motion studies. To our best knowledge, this is the first large-scale video dataset recorded on factory floors and labeled by industrial engineers. Below we first provide some details of our IndustryASM dataset and then describe how we convert timestamp and action labels into chat samples.
4.1 Dataset Details
Our IndustryASM dataset includes 4803 videos with each video lasting 119 seconds on average, yielding 158 hours as the total dataset duration. In addition, it consists of 52 various industry assembly procedures with each assembly procedure having 12 steps on average. Example videos are shown in Fig. 3. Due to space limits, we provide additional details of IndustryASM in the supplementary material.
4.2 Chat Samples
Our industry assembly videos are manually annotated by industrial engineers into relevant framewise action segments. The action labels contain the same step naming conventions as used on factory floors during manufacturing processes. Given the framewise annotations we first convert them to timestamps by using the frame rate of the video. We manually write the instructions designed specifically for the action segmentation task to ensure quality. Then we generate the answers by using the timestamps and the action labels.

### v2

@@ Abstract
This paper introduces TemporalVLM, a video large language model (video LLM) capable of effective temporal reasoning and fine-grained understanding in long videos. At the core, our approach includes a visual encoder for mapping a long-term input video into features which are time-aware and contain both local and global cues. In particular, it first divides the input video into short-term clips, which are jointly encoded with their timestamps into time-sensitive local features. Next, the local features are passed through a bidirectional long short-term memory (BiLSTM) module for global feature aggregation. The extracted time-aware and multi-level features are important for accurate temporal reasoning and fine-grained understanding in long videos. Moreover, to facilitate the evaluation of TemporalVLM, we present a large-scale long video dataset of industry assembly processes, namely IndustryASM, which consists of videos recorded on factory floors with actions and timestamps annotated by industrial engineers for time and motion studies and temporal action segmentation evaluation. Finally, extensive experiments on datasets of long videos, including TimeIT and IndustryASM, show that TemporalVLM achieves superior performance than previous methods across temporal reasoning and fine-grained understanding tasks, namely dense video captioning, temporal video grounding, video highlight detection, and temporal action segmentation. To the best of our knowledge, our work is the first to incorporate LSTMs into video LLMs.
@@ 1 Introduction
Video temporal reasoning represents the process of reasoning about time and its passage in videos, with a focus on how events or actions happen and relate to each other in terms of time. Several video understanding applications require the ability of temporal reasoning. Examples include dense video captioning [45, 59, 52], which attempts to detect the timestamps and generate detailed descriptions of all the events or actions in an input video, and temporal action segmentation [22, 53, 2, 25], which aims to associate each frame of a long video capturing a complex activity with one of the action/sub-activity classes. Considerable efforts [52, 26, 29, 25] have been invested in developing separate models for solving individual tasks. These models usually have different architectures. Therefore, it is favorable to design a unified model for handling various tasks.
@@
In this paper, we propose TemporalVLM, a video LLM capable of effective temporal reasoning and fine-grained understanding in long videos. Fig. 1 shows the differences in model architectures of TemporalVLM and previous works. Our model includes two key architectural contributions for temporal reasoning in long videos. Firstly, we divide a long-term input video into multiple short-term clips and introduce a time-sensitive clip encoder for extracting time-aware local features from each clip. Secondly, we adopt a bidirectional long short-term memory (BiLSTM) module which takes all the local features as inputs and computes global features from multiple clips. As a consequence, our features not only are time-sensitive but also contain both local fine-grained and global semantic information, which are crucial for temporal reasoning in long videos. Moreover, to further evaluate our model, we present IndustryASM, a large-scale long video dataset of industry assembly processes for temporal action segmentation benchmarking and time and motion studies. Our IndustryASM dataset comprises of $4803$ videos with an average video duration of $119$ seconds. It covers in total $52$ diverse industry assembly tasks and includes timestamp and action labels. We convert the labels into chat samples with manually written instructions. To the best of our knowledge, it is the first large-scale video dataset recorded on factory floors and labeled by industrial engineers for temporal action segmentation evaluation and time and motion analyses. Lastly, extensive experiments on datasets of long videos, i.e., TimeIT [35] and IndustryASM, demonstrate that our TemporalVLM approach achieves superior results over previous methods across temporal reasoning and fine-grained understanding tasks, namely dense video captioning, temporal video grounding, video highlight detection, and temporal action segmentation.
@@
  • We develop TemporalVLM, a video LLM for temporal reasoning and fine-grained understanding in long videos. Given a long-term input video, we first sample short-term clips and encode each clip along with its timestamps into time-aware local features. The local features are then fed to a BiLSTM module to extract global features. Therefore, our features not only are time-sensitive but also capture effective local and global information.
@@
  • Our evaluations show that TemporalVLM outperforms prior works across temporal reasoning and fine-grained understanding tasks. To our best knowledge, this is the first work to integrate LSTMs into video LLMs.
@@ 2 Related Work
Video Large Language Models. Motivated by the impressive reasoning and generalization abilities of LLMs, video LLMs [21, 23, 27, 28, 56, 15, 38] have been developed for video understanding tasks. They typically include three key components: i) a pre-trained visual encoder (e.g., ViT [37]) to extract visual features, ii) a projection layer to map visual features into the text latent space of LLMs, and iii) a pre-trained LLM (e.g., LLaMA [41]) for generating responses. They mostly differ in their visual encoder, i.e., how visual features are extracted and aggregated from an input video. For example, VideoChat [21] extracts frame features via a visual transformer [37] and employs a query transformer (Q-Former) [20] to aggregate frame features into video features, while a video Q-Former is further included for temporal modeling in Video-LLaMA [56]. The above methods usually map the video into a fixed number of visual tokens, yielding degrading performance with long videos, while encoding video frames and their timestamps separately and hence struggling with temporal reasoning tasks. To address these drawbacks, methods with temporal reasoning capabilities have been introduced, e.g., TimeChat [35] presents a sliding video Q-Former to handle various video lengths and a time-aware frame encoder to jointly encode video frames and their timestamps. These methods often consider the video as a single clip [27, 28, 21, 56] and aggregate video tokens via pooling operation [27, 28] and query aggregation [21, 56], overlooking fine-grained details. In this work, we divide the video into multiple clips and propose a time-aware clip encoder for capturing local features from each clip. Also, we introduce a BiLSTM module for aggregating global features over multiple clips. To our best knowledge, our work is the first to exploit LSTMs for video LLMs.
@@
Long Video Understanding. There exist a few challenges in long video understanding such as complex spatial-temporal relationships and redundant information. Several methods have been developed for long video understanding, including efficient architectures [6, 17], temporal pooling/aggregation [36, 50], and dynamic clip selection [9, 10]. For vision-language understanding tasks, methods based on temporal alignment [11, 3] have been introduced, while memory techniques [51, 57] are often used in dense video prediction tasks. Recently, Weng et al. [49] propose LongVLM for long video understanding. It divides the video into clips and employs a merging module for extracting local features and pooling operation for computing global features from multiple clips. It does not utilize timestamps, which are crucial for temporal reasoning. In contrast, our TemporalVLM model explicitly utilizes timestamps via a time-aware clip encoder, yielding time-aware features. Moreover, we employ a learnable BiLSTM module for aggregating global features from multiple clips.
@@ 3 TemporalVLM: A Video LLM for Temporal Reasoning in Long Videos
Figure 2: Given a long-term input video with timestamps, we divide it into short-term clips with timestamps. The time-aware clip encoder is then employed on each clip along with its timestamps for extracting time-aware local features, which capture local fine-grained cues within each clip. Next, the bidirectional long short-term memory module is adopted for computing global features, which capture global temporal relationships across multiple clips. Lastly, the large language model takes video and text tokens as inputs and generates responses.
Below we describe our first contribution, TemporalVLM, a video LLM for effective temporal reasoning in long videos. Two novel components are proposed in TemporalVLM: i) a time-aware clip encoder for extracting time-aware local features from a short-term clip, and ii) a BiLSTM module for computing global features from the local features. We provide an overview of our TemporalVLM model in Fig. 2.
@@ 3.1 Time-Aware Clip Encoder
Video LLMs often consider a long-term input video as a single clip and uniformly sample a fixed number of frames $N^{v}_{f}$ from the video, e.g., $N^{v}_{f}=96$ frames in TimeChat [35] and $N^{v}_{f}=100$ frames in VTimeLLM [13]. Moreover, prior works usually use query aggregation, e.g., TimeChat [35], and pooling operation, e.g., Video-ChatGPT [28], to aggregate video tokens. Thus, they often perform well with short-term videos but miss out important fine-grained information for long-term videos, leading to degrading performance.
We propose a time-aware clip encoder (i.e., blue block in Fig. 2) for extracting local fine-grained cues within a clip. Given a long-term input video, we first decompose it into $C$ short-term clips and uniformly sample a fixed number of frames $N^{c}_{f}$ from each clip. Here, we set $C=6$ clips and $N^{c}_{f}=96$ frames per clip. The sampled frames for each clip along with their corresponding timestamps are then passed to our time-aware clip encoder for simultaneously encoding the frame contents and their timestamps, yielding time-aware local features. We adopt the time-aware frame encoder and sliding video Q-Former of TimeChat [35] as our time-aware clip encoder. The time-aware frame encoder uses a pre-trained image encoder [39] to obtain frame features, which are jointly encoded with timestamps via an image Q-Former [5], yielding time-aware frame features. These frame features are then fed to a sliding video Q-Former [35] for temporal modeling, yielding time-aware clip features for each clip. Recent works, e.g., LongVLM [49], leverage clip sampler and clip encoder for long video understanding but their encoders are not time-aware and hence are not effective for temporal reasoning.
@@ 3.2 Bidirectional Long Short-Term Memory
Our time-sensitive clip encoder is applied on each short-term clip to obtain time-aware local features, which capture useful fine-grained cues for temporal reasoning within the short-term clip. However, they are not effective for temporal reasoning over the long-term video, which requires the ability to capture long-range temporal dependencies.
We introduce a BiLSTM module (i.e., green block in Fig. 2) for computing global features across multiple clips. We first concatenate time-aware local features extracted from clips in the temporal order which the clips appear in the video. We then pass the sequence of local features to a BiLSTM module for aggregating global features. Our BiLSTM module follows a standard architecture, including two LSTM networks: one processes the sequence in the original order (forward LSTM) and another processes the sequence in the reverse order (backward LSTM) as:
$\mathbf{h}^{f}_{t}=\text{LSTM}(\mathbf{h}^{f}_{t-1},\mathbf{c}_{t}),$ (1)
$\mathbf{h}^{b}_{t}=\text{LSTM}(\mathbf{h}^{b}_{t+1},\mathbf{c}_{t}),$ (2)
where $\mathbf{h}^{f}_{t}$ and $\mathbf{h}^{b}_{t}$ denote the hidden states at time step $t$ of the forward LSTM and backward LSTM respectively and $\mathbf{c}_{t}$ is the input at time step $t$. The final output $\mathbf{h}_{t}$ at time step $t$ is obtained by concatenating the outputs of the forward LSTM and backward LSTM as:
$\mathbf{h}_{t}=[\mathbf{h}^{f}_{t},\mathbf{h}^{b}_{t}].$ (3)
Please refer to the textbook [18] for a detailed description. Our BiLSTM module utilizes both past and future information and is capable of capturing long-range temporal relationships in both forward and backward directions. As empirically observed in Sec. 5.3, BiLSTM outperforms various alternatives, including average pooling, linear layer, traditional LSTM, and transformer [43]. Our final features output by BiLSTM not only are time-sensitive but also contain both local fine-grained and global semantic information, which are important for temporal reasoning in long videos. Unlike the recent LongVLM model [49], which applies pooling operation to obtain global features, our BiLSTM module has learnable parameters and achieves better results as shown in Sec. 5.
The video tokens output by our BiLSTM module have dimensions of $(C*N^{c}_{f},~{}2*N_{V})$, while the LLM requires an input of dimensions $(N^{c}_{f},~{}N_{LLM})$. Here, $N_{V}$ and $N_{LLM}$ are the dimensions of the video tokens and the LLM latent space respectively, and $2$ is to account for the outputs of both forward and backward passes of BiLSTM. Thus, we pass the video tokens output by BiLSTM through projection layers to match the input dimensions required by the LLM.
@@ 3.3 Large Language Model
The LLM (i.e., yellow block in Fig. 2) takes as input the video tokens $\mathbf{X}^{v}$, query tokens $\mathbf{X}^{q}$, and generates responses $\mathbf{X}^{r}$ to users. Video LLMs typically employ a two-stage training scheme: i) leveraging large-scale image/video-text pairs for vision-language alignment to pre-train the model, and ii) utilizing instruction data for instruction following to fine-tune the pre-trained model. Following TimeChat [35], we use the checkpoint of the open-source LLaMA-2 7B model [42] and perform instruction tuning only. We optimize the below loss function:
$\mathcal{L}=-\log P_{\mathbf{\psi}}(\mathbf{X}^{r}|\mathbf{X}^{v},\mathbf{X}^{q})$ (4)
$=-\sum^{L}_{i=1}\log P_{\mathbf{\psi}}(x_{i}|\mathbf{X}^{v},\mathbf{X}^{q},\mathbf{X}^{r}_{<i}).$ (5)
Here, $\mathbf{\psi}$ is the learnable parameters of TemporalVLM, $L$ is the response length, $x_{i}$ is the current predicted token, and $\mathbf{X}^{r}_{<i}$ is the prior tokens appearing before $x_{i}$ in the response. Also, we employ LoRA [12] for efficient fine-tuning.
@@ 4 IndustryASM: A Large-Scale Long Video Dataset of Industry Assembly Processes
We follow prior works [32, 13, 14, 35] to test TemporalVLM on dense video captioning, temporal video grounding, and video highlight detection. To further evaluate our model, we introduce IndustryASM, a large-scale long video dataset of industry assembly processes for temporal action segmentation and time and motion studies. To our best knowledge, it is the first large-scale video dataset recorded on factory floors and labeled by industrial engineers. Below we provide some details and annotation process.
4.1 Dataset Statistics
Our IndustryASM dataset includes 4803 videos with each video lasting 119 seconds on average, yielding 158 hours as the total dataset duration. In addition, it consists of 52 various industry assembly procedures with each assembly procedure having 12 steps on average. Example videos are shown in Fig. 3. Due to space limits, we provide additional details of IndustryASM in our supplementary material.
4.2 Annotation Process
Our industry assembly videos are manually annotated by industrial engineers into relevant framewise action segments. The action labels contain the same step naming conventions as used on factory floors during manufacturing processes. To ensure quality, each video is annotated by two labelers, i.e., one labeler obtains the labels for the video, while another labeler checks the labels. In cases of conflicts, both labelers discuss to address them. Overall, around 8% of the videos have conflicts and require fixing, yielding an agreement rate of roughly 92% between labelers. Given the framewise annotations we first convert them to timestamps by using the frame rate of the video. We manually write the instructions designed for the action segmentation task to ensure quality and generate the answers by using timestamps and action labels. Example instructions and responses are shown in Fig. S1 of our supplementary material.